A regular hexagon is inscribed in a circle with 2 inch radius .what is the side length of a square with the same area as the hexagon.

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
Did you get 3sqrt3 as the area of the hexagon?
-----
If you did your next step is:
s^2=3sqrt3, where s^2 is the area of a square with side = s.
Then s^2 = 3^(3/2) 
Then s = 3^(3/4) (final answer) 
cheers,
 
Regards
Arun (askIITians forum expert)